# VISTA: Enhancing Vision-Text Alignment in MLLMs via Cross-Modal Mutual Information Maximization

### v1

@@ 1 Introduction
The rapid advancement of artificial intelligence has propelled expectations for AI assistants beyond traditional text-based interactions, driving demand for systems that align more closely with the inherently multimodal nature of human perception and communication. Multimodal Large Language Models (MLLMs) have emerged at the forefront of this pursuit, with architectures like LLaVA [1] pioneering a widely adopted paradigm. These models typically utilize a Multi-Layer Perceptron (MLP) to bridge pre-trained Vision Transformers (ViTs) and Large Language Models (LLMs), integrating visual and textual modalities. This “Glue Structure” approach, exemplified by influential models such as InternVL [2], QwenVL [3], and Kimi-VL [4], facilitates rapid, cost-effective adaptation of the latest unimodal foundation models, enabling the swift development of new MLLMs and transcending the limitations of text-only systems to address complex multimodal problems.
@@
Despite their architectural adaptability, prevalent MLLMs predominantly rely on standard cross-entropy loss applied to text generation for end-to-end training and alignment. This straightforward approach inherently prioritizes the accurate prediction of text tokens. Consequently, it often leads to two critical issues: (1) a bias towards textual information, potentially diminishing the significance of visual context in semantic understanding, and (2) an implicit and often weak vision-text alignment mechanism in text-only cross-entropy loss that lacks explicit constraints to enforce semantic consistency between visual and textual representations. Recent studies [5, 6, 7] corroborate this, highlighting that existing MLLMs frequently exhibit a modality imbalance, with their representational capacity heavily skewed towards textual information.
@@
Several attempts have been made to mitigate this imbalance by incorporating forms of image-focused supervision. For instance, ROSS [8] introduces auxiliary learnable modules to reconstruct input images during alignment training, thereby enhancing fine-grained visual perception. Similarly, Metamorph [9] appends a trainable MLP to reconstruct visual semantic vectors from the ViT, aiming to create a unified MLLM for both generation and comprehension. While these methods have shown improvements, they often introduce additional trainable parameters, require architectural modifications, or necessitate extra training data and stages, and may not directly address the core issue of implicit cross-modal semantic alignment.
@@
The limitations of current implicit alignment paradigms are starkly illustrated in Figure 1, where visualizations of image-text embedding similarity in models like LLaVA reveal considerable semantic noise between modalities. This underscores a fundamental weakness: the absence of a direct, explicit objective to align visual and textual representations in a shared semantic space. In contrast, foundational models like CLIP [10] have demonstrated the power of explicit alignment by directly optimizing the correspondence between image and text embeddings. Recent work [11] further reinforces the potential of explicit alignment strategies by employing masked autoencoder-style post-training on ViTs to enhance local-region awareness within the visual-textual semantic space. This suggests that incorporating explicit alignment objectives is a promising avenue for overcoming the current limitations of MLLMs.
@@
Figure 1: Visualization of the semantic similarity between prompt text embeddings and image embeddings generated by the MLP. The prompt used is: "How do you solve this specific type of puzzle cube depicted in the image?" Similarity is measured using cosine similarity with mean aggregation. Darker colors indicate higher semantic alignment.
@@ 2 Related work
Architectures of multimodal large language models. The design of multimodal large language models (MLLMs) has evolved from early query-based cross-modal interaction frameworks, exemplified by MiniGPT-4 [12] and BLIP-2 [13], which utilize modules like Q-Former to mediate between visual and textual modalities. A paradigm shift occurred with LLaVA [1], introducing a streamlined encoder-MLP-LLM pipeline that projects frozen vision encoder (e.g., ViT [14]) features into the LLM embedding space via a lightweight MLP. This architecture has become the prevailing standard, adopted by subsequent models such as InternVL [2], QwenVL [3, 15], and KimiVL [4]. Although these methods input image tokens as prompts, they still optimize only the LLM’s text-only cross-entropy loss. Thus, from the decoder’s perspective, alignment remains implicit, with no explicit supervision linking image tokens to text.
Challenges in vision-language alignment. Recent works reveal fundamental limitations in current alignment strategies, particularly their failure to preserve detailed visual semantics within MLLMs. For instance, InternVL-1.5 [2] showed that scaling vision encoders from scratch significantly improves visual grounding, while InternVL-3 [16] enhanced semantic integration by co-training ViT and LLM parameters. Auxiliary modules like Metamorph [9], AIMV2 [17], and Ross [8] further encourage visual concept internalization or semantic consistency via MLP reconstruction, pixel alignment, or diffusion-based modules. Although empirically effective, these methods rely on task-specific modules, extensive end-to-end training, or architectural changes, lacking a unified theoretical framework and often overlooking the mismatch between MLLMs’ instruction-following objectives and the semantic fidelity needed for robust visual understanding.
Explicit alignment in multimodal pretraining. Encoder-level explicit alignment has shown effectiveness in multimodal representation learning, exemplified by CLIP [10], which maximizes mutual information between image-text embeddings, and GRAM [18], which enforces geometric alignment by minimizing the Gram volume spanned by modality vectors, achieving state-of-the-art results on video-audio-text benchmarks. In contrast, MLLMs mainly depend on implicit alignment via autoregressive text generation, which remains insufficient to bridge the modality gap [5, 19, 20]. Recent attempts to incorporate explicit alignment into MLLMs include Sugar [21], introducing a trainable MLP and interlaced image-text data but altering training data format, and a masked joint training approach [11] optimizing encoder-level alignment through the text encoder. Despite progress, a principled, generalizable explicit alignment framework tailored for MLLMs remains an open challenge.
@@ 3.2 Vanish of visual contribution as the text length grows
Lemma 1 (Autoregressive entropy growth)
@@
Proof 1
By the chain rule of entropy and non-degeneracy assumption, each conditional entropy term is lower bounded by $\Delta_{H}$. Summing over $t-1$ terms yields the linear growth.
Lemma 2 (Bounded visual mutual information)
@@
Proof 2
By definition, $\mathcal{I}(x^{T}_{t};\mathbf{S}^{I})=\mathcal{H}(x^{T}_{t})-\mathcal{H}(x^{T}_{t}\mid\mathbf{S}^{I})\leq\mathcal{H}(x^{T}_{t})$. Since $x^{T}_{t}$ is discrete with bounded vocabulary, its entropy is upper bounded by $C$.
Theorem 1 (Vanishing vision-text alignment contribution)
@@
Proof 3
From Lemma 1, the textual mutual information grows at least linearly with $t$. From Lemma 2, the visual mutual information is upper bounded by $C$. Since:
@@
and $\mathcal{I}(x^{T}_{t};s^{T}_{<t}\mid\mathbf{S}^{I})$ grows unboundedly with $t$, the ratio tends to zero, see Appendix A.
@@ 4.1 Enhancing cross-modal alignment via explicit regularization
Building upon the theoretical foundation established in Theorem 1, we identify a fundamental limitation inherent to the implicit alignment objective: as the proportion of textual tokens grows during training, the strength of vision-text alignment progressively weakens, thereby impairing effective cross-modal integration. To mitigate this issue, we introduce an explicit alignment regularizer that directly enforces stronger correspondence between visual and textual representations.
@@ 5.1 Experimental setup
Datasets. To ensure a fair and rigorous evaluation, we follow the dataset configuration used in LLaVA-v1.5. For pre-training, a subset of the LAION/CC/SBU dataset filtered for balanced concept coverage and augmented with BLIP-generated captions is employed, consistent with LLaVA’s setup. Instruction tuning utilizes a combination of COCO [22], GQA [23], OCR-VQA [24], TextVQA [25], and VisualGenome [26], as in [27]. Further dataset details can be found therein.
Tasks and evaluation. Building upon prior benchmarks, we conduct comprehensive evaluations across three categories: (1) high-level semantic VQA tasks, encompassing VQAv2 [28], OK-VQA [29], GQA [23], TextVQA [25], RealWorldQA [30], and DocVQA [31]; (2) general multimodal understanding benchmarks covering diverse scenarios, including MMBench [32], SEED [33], AI2D [34], MMMU [35], MMTB [36], OCR-VQA [24], and MMStar [37]; and (3) fine-grained visual perception and retrieval tasks, evaluated on MME [38] alongside RefCOCO [39], RefCOCO+ [39], and RefCOCOg [40]. All evaluations are conducted using the lmms-eval framework [41, 42], as recommended by LLaVA, which integrates multiple metrics to provide a unified and thorough assessment. Detailed dataset and metric descriptions are provided in Appendix C.
Models. To rigorously validate VISTA, we experiment with two MLLM backbones of distinct scales: TinyLLaVA [43] (3B parameters) and LLaVA-v1.5 [27] (7B parameters). Both baselines are trained and fine-tuned on identical datasets to ensure comparability. For TinyLLaVA, we adopt the phi-2 [44] 3B model as the language backbone and SigLIP [45] as the vision encoder. This setup enables a thorough assessment of VISTA’s generalizability and efficacy across varying model capacities and architectures.
Baselines and implementation. Following the data and training protocols in [27], we compare TinyLLaVA-VISTA with TinyLLaVA, and LLaVA-v1.5-VISTA with LLaVA-v1.5. The same weighting function $f(t)$ is used for both scales without hyperparameter tuning. Detailed model and training settings are provided in the appendix. During pre-training and fine-tuning, we optimize a combined loss of VISTA alignment and standard cross-entropy.
@@ 5.3 Case study
Figure 3: Semantic similarity visualization between MLP-generated prompt embeddings and image embeddings on LLaVA Wild Bench[1]. Each subfigure presents two images per case: left—LLaVA baseline; right—our VISTA method. Darker colors indicate higher semantic alignment.
@@ 6 Conclusion
In this work, we tackled the challenge of suboptimal vision-language alignment in MLLMs, often caused by limitations of conventional text-based cross-entropy loss. Our information-theoretic analysis showed that this implicit alignment weakens as text length grows, reducing effective visual integration. To address this, we propose VISTA, a simple and efficient method that explicitly maximizes cross-modal mutual information without adding parameters, changing architectures, or requiring extra data. Extensive experiments across diverse benchmarks demonstrate VISTA’s significant improvements in visual understanding and reasoning.
These results highlight the value of explicit alignment objectives and position VISTA as a practical, scalable solution for building more capable and balanced MLLMs, paving the way for future research in principled cross-modal alignment.
@@ Appendix A Detailed Proofs
A.1 Proof of Lemma 1 (Autoregressive Entropy Growth)
Proof 4
@@ A.1 Proof of Lemma 1 (Autoregressive Entropy Growth)
A.2 Proof of Lemma 2 (Bounded Visual Mutual Information)
Proof 5
@@ A.2 Proof of Lemma 2 (Bounded Visual Mutual Information)
A.3 Proof of Theorem 1 (Vanishing Visual Alignment Contribution)
Proof 6
@@ A.3 Proof of Theorem 1 (Vanishing Visual Alignment Contribution)
First, from Lemma 2, the mutual information between the visual signals and any single token is upper bounded by a constant:
@@
From Lemma 1, the mutual information between the current token and the preceding text context grows at least linearly with $t$:
@@ C.1 Visual question answering benchmarks
OK-VQA: OK-VQA(Outside Knowledge VQA)[29] is a visual question answering dataset that requires external knowledge. The answers to the questions cannot be inferred solely from the image but also need to incorporate common sense or world knowledge. This dataset evaluates the model’s ability in the intersection of vision and knowledge reasoning. We used the acc of validation set to report Our result.
GQA: GQA(Graph Question Answering)[23] generates questions and answers based on image scene graphs, focusing on structured reasoning. It emphasizes logical analysis and challenges the model’s depth of understanding of semantics and context.
VQAV2: VQAv2[28] is a new dataset containing open-ended questions about images. These questions require an understanding of vision, language and commonsense knowledge to answer. We used the test split to report Our result.
TextVQA: TextVQA[25] focuses on textual information in images, requiring models to recognize and comprehend text within images to answer questions. It drives research on the integration of visual and textual information, expanding the boundaries of visual question answering.
RealWorldQA: RealWorldQA[30] features images and questions sourced from real-world scenarios, encompassing diverse content from daily life. The dataset imposes higher requirements on the model’s generalization ability and adaptability to complex scenes.
DOCVQA: This dataset[31] comprises over 50,000 questions formulated on more than 12,000 document images. Document Visual Question Answering (DocVQA) aims to promote a "purpose-driven" perspective within the field of Document Analysis and Recognition, encouraging research that directly addresses practical and semantic understanding of document content.

### v2

@@ 1 Introduction
The rapid advancement of artificial intelligence has propelled expectations for AI assistants beyond traditional text-based interactions, driving demand for systems that align more closely with the inherently multimodal nature of human perception and communication. Multimodal Large Language Models (MLLMs) have emerged at the forefront of this pursuit, with architectures like LLaVA (Liu et al., 2024b) pioneering a widely adopted paradigm. These models typically utilize a Multi-Layer Perceptron (MLP) to bridge pre-trained Vision Transformers (ViTs) and Large Language Models (LLMs), integrating visual and textual modalities. This “Glue Structure” approach, exemplified by influential models such as InternVL (Chen et al., 2024b), QwenVL (Wang et al., 2024b), and Kimi-VL (Team et al., 2025), facilitates rapid, cost-effective adaptation of the latest unimodal foundation models, enabling the swift development of new MLLMs and transcending the limitations of text-only systems to address complex multimodal problems.
@@
Despite their architectural adaptability, prevalent MLLMs predominantly rely on standard cross-entropy loss applied to text generation for end-to-end training and alignment. This straightforward approach inherently prioritizes the accurate prediction of text tokens. Consequently, it often leads to two critical issues: (1) a bias towards textual information, potentially diminishing the significance of visual context in semantic understanding, and (2) an implicit and often weak vision-text alignment mechanism in text-only cross-entropy loss that lacks explicit constraints to enforce semantic consistency between visual and textual representations. Recent studies (Huang et al., 2024; Zhang et al., 2024b; Liu et al., 2024c) corroborate this, highlighting that existing MLLMs frequently exhibit a modality imbalance, with their representational capacity heavily skewed towards textual information.
@@
Several attempts have been made to mitigate this imbalance by incorporating forms of image-focused supervision. For instance, ROSS (Wang et al., 2024a) introduces auxiliary learnable modules to reconstruct input images during alignment training, thereby enhancing fine-grained visual perception. Similarly, Metamorph (Tong et al., 2024) appends a trainable MLP to reconstruct visual semantic vectors from the ViT, aiming to create a unified MLLM for both generation and comprehension. While these methods have shown improvements, they often introduce additional trainable parameters, require architectural modifications, or necessitate extra training data and stages, and may not directly address the core issue of implicit cross-modal semantic alignment.
@@
The limitations of current implicit alignment paradigms are starkly illustrated in Figure 2, where visualizations of image-text embedding similarity in models like LLaVA reveal considerable semantic noise between modalities. This underscores a fundamental weakness: the absence of a direct, explicit objective to align visual and textual representations in a shared semantic space. In contrast, foundational models like CLIP (Radford et al., 2021) have demonstrated the power of explicit alignment by directly optimizing the correspondence between image and text embeddings. Recent work (Covert et al., 2024) further reinforces the potential of explicit alignment strategies by employing masked autoencoder-style post-training on ViTs to enhance local-region awareness within the visual-textual semantic space. This suggests that incorporating explicit alignment objectives is a promising avenue for overcoming the current limitations of MLLMs.
@@
Figure 2: Visualization of the semantic similarity between prompt text embeddings and image embeddings generated by the MLP. The prompt used is: ”How do you solve this specific type of puzzle cube depicted in the image?” Similarity is measured using cosine similarity with mean aggregation. Darker colors indicate higher semantic alignment.
@@ 2 Related work
Architectures of multimodal large language models. The design of multimodal large language models (MLLMs) has evolved from early query-based cross-modal interaction frameworks, exemplified by MiniGPT-4 Zhu et al. (2023) and BLIP-2 Li et al. (2023b), which utilize modules like Q-Former to mediate between visual and textual modalities. A paradigm shift occurred with LLaVA Liu et al. (2024b), introducing a streamlined encoder-MLP-LLM pipeline that projects frozen vision encoder (e.g., ViT Dosovitskiy et al. (2020)) features into the LLM embedding space via a lightweight MLP. This architecture has become the prevailing standard, adopted by subsequent models such as InternVL Chen et al. (2024b), QwenVL Wang et al. (2024b); Bai et al. (2025), and KimiVL Team et al. (2025). Although these methods input image tokens as prompts, they still optimize only the LLM’s text-only cross-entropy loss. Thus, from the decoder’s perspective, alignment remains implicit, with no explicit supervision linking image tokens to text.
Challenges in vision-language alignment. Recent works reveal fundamental limitations in current alignment strategies, particularly their failure to preserve detailed visual semantics within MLLMs. For instance, InternVL-1.5 Chen et al. (2024b) showed that scaling vision encoders from scratch significantly improves visual grounding, while InternVL-3 Zhu et al. (2025) enhanced semantic integration by co-training ViT and LLM parameters. Auxiliary modules like Metamorph Tong et al. (2024), AIMV2 Fini et al. (2024), and Ross Wang et al. (2024a) further encourage visual concept internalization or semantic consistency via MLP reconstruction, pixel alignment, or diffusion-based modules. Although empirically effective, these methods rely on task-specific modules, extensive end-to-end training, or architectural changes, lacking a unified theoretical framework and often overlooking the mismatch between MLLMs’ instruction-following objectives and the semantic fidelity needed for robust visual understanding.
Explicit alignment in multimodal pretraining. Encoder-level explicit alignment has shown effectiveness in multimodal representation learning, exemplified by CLIP Radford et al. (2021), which maximizes mutual information between image-text embeddings, and GRAM Cicchetti et al. (2024), which enforces geometric alignment by minimizing the Gram volume spanned by modality vectors, achieving state-of-the-art results on video-audio-text benchmarks. In contrast, MLLMs mainly depend on implicit alignment via autoregressive text generation, which remains insufficient to bridge the modality gap Huang et al. (2024); Wu et al. (2025); Li et al. (2025). Recent attempts to incorporate explicit alignment into MLLMs include Sugar Chow et al. (2024), introducing a trainable MLP and interlaced image-text data but altering training data format, and a masked joint training approach (Covert et al., 2024) optimizing encoder-level alignment through the text encoder. Despite progress, a principled, generalizable explicit alignment framework tailored for MLLMs remains an open challenge.
@@ 3.2 Vanish of visual contribution as the text length grows
Lemma 3.1 (Autoregressive entropy growth).
@@
Proof.
By the chain rule of entropy and non-degeneracy assumption, each conditional entropy term is lower bounded by $\Delta_{H}$. Summing over $t-1$ terms yields the linear growth. ∎
Lemma 3.2 (Bounded visual mutual information).
@@
Proof.
By definition, $\mathcal{I}(x^{T}_{t};\mathbf{S}^{I})=\mathcal{H}(x^{T}_{t})-\mathcal{H}(x^{T}_{t}\mid\mathbf{S}^{I})\leq\mathcal{H}(x^{T}_{t})$. Since $x^{T}_{t}$ is discrete with bounded vocabulary, its entropy is upper bounded by $C$. ∎
Theorem 3.3 (Vanishing vision-text alignment contribution).
@@
Proof.
From Lemma 3.1, the textual mutual information grows at least linearly with $t$. From Lemma 3.2, the visual mutual information is upper bounded by $C$. Since:
@@
and $\mathcal{I}(x^{T}_{t};s^{T}_{<t}\mid\mathbf{S}^{I})$ grows unboundedly with $t$, the ratio tends to zero, see Appendix A. ∎
@@ 4.1 Enhancing cross-modal alignment via explicit regularization
Building upon the theoretical foundation established in Theorem 3.3, we identify a fundamental limitation inherent to the implicit alignment objective: as the proportion of textual tokens grows during training, the strength of vision-text alignment progressively weakens, thereby impairing effective cross-modal integration. To mitigate this issue, we introduce an explicit alignment regularizer that directly enforces stronger correspondence between visual and textual representations.
@@ 5.1 Experimental setup
Datasets. To ensure a fair and rigorous evaluation, we follow the dataset configuration used in LLaVA-v1.5. For pre-training, a subset of the LAION/CC/SBU dataset filtered for balanced concept coverage and augmented with BLIP-generated captions is employed, consistent with LLaVA’s setup. Instruction tuning utilizes a combination of COCO Lin et al. (2014), GQA Hudson and Manning (2019), OCR-VQA Mishra et al. (2019), TextVQA Singh et al. (2019), and VisualGenome Krishna et al. (2017), as in Liu et al. (2024a). Further dataset details can be found therein.
Tasks and evaluation. Building upon prior benchmarks, we conduct comprehensive evaluations across three categories: (1) high-level semantic VQA tasks, encompassing VQAv2 Goyal et al. (2017), OK-VQA Marino et al. (2019), GQA Hudson and Manning (2019), TextVQA Singh et al. (2019), RealWorldQA x.ai (2024), and DocVQA Mathew et al. (2021); (2) general multimodal understanding benchmarks covering diverse scenarios, including MMBench Liu et al. (2025), SEED Li et al. (2023a), AI2D Kembhavi et al. (2016), MMMU Yue et al. (2024), MMTB Ying et al. (2024), OCR-VQA Mishra et al. (2019), and MMStar Chen et al. (2024a); and (3) fine-grained visual perception and retrieval tasks, evaluated on MME Fu et al. (2024) alongside RefCOCO Kazemzadeh et al. (2014), RefCOCO+ Kazemzadeh et al. (2014), and RefCOCOg Mao et al. (2016). All evaluations are conducted using the lmms-eval framework Zhang et al. (2024a); Li et al. (2024), as recommended by LLaVA, which integrates multiple metrics to provide a unified and thorough assessment. Detailed dataset and metric descriptions are provided in Appendix C.
Models. To rigorously validate VISTA, we experiment with two MLLM backbones of distinct scales: TinyLLaVA Zhou et al. (2024) (3B parameters) and LLaVA-v1.5 Liu et al. (2024a) (7B parameters). Both baselines are trained and fine-tuned on identical datasets to ensure comparability. For TinyLLaVA, we adopt the phi-2 Javaheripi et al. (2023) 3B model as the language backbone and SigLIP Zhai et al. (2023) as the vision encoder. This setup enables a thorough assessment of VISTA’s generalizability and efficacy across varying model capacities and architectures.
Baselines and implementation. Following the data and training protocols in Liu et al. (2024a), we compare TinyLLaVA-VISTA with TinyLLaVA, and LLaVA-v1.5-VISTA with LLaVA-v1.5. The same weighting function $f(t)$ is used for both scales without hyperparameter tuning. Detailed model and training settings are provided in the appendix. During pre-training and fine-tuning, we optimize a combined loss of VISTA alignment and standard cross-entropy.
@@ 5.3 Case study
Figure 3: Semantic similarity visualization between MLP-generated prompt embeddings and image embeddings on LLaVA Wild BenchLiu et al. (2024b). Each subfigure presents two images per case: left—LLaVA baseline; right—our VISTA method. Darker colors indicate higher semantic alignment.
@@ 6 Conclusion
In this work, we addressed the critical challenge of suboptimal vision-language alignment in MLLMs, which often stems from the limitations of conventional text-based cross-entropy loss. Our information-theoretic analysis revealed that the implicit alignment fostered by this loss degrades as text length increases, diminishing the effective integration of visual information. In response to this problem, we propose a novel method VISTA that introduces an explicit objective to maximize cross-modal mutual information. A key advantage of VISTA is its simplicity and efficiency, enhancing alignment without necessitating additional parameters, architectural changes, or supplementary training data.
Our extensive experimental results across a wide array of benchmarks robustly demonstrate VISTA’s effectiveness, yielding significant performance improvements in visual understanding and reasoning tasks. These findings underscore the benefits of explicit alignment objectives and suggest that VISTA offers a practical and scalable approach to building more capable and balanced MLLMs. We believe this work not only provides a valuable tool for current MLLMs development but also opens promising avenues for future research into principled and efficient cross-modal alignment techniques.
@@ Appendix A Detailed Proofs
A.1 Proof of Lemma 3.1 (Autoregressive Entropy Growth)
Proof.
@@
A.2 Proof of Lemma 3.2 (Bounded Visual Mutual Information)
Proof.
@@
A.3 Proof of Theorem 3.3 (Vanishing Visual Alignment Contribution)
Proof.
@@
First, from Lemma 3.2, the mutual information between the visual signals and any single token is upper bounded by a constant:
@@
From Lemma 3.1, the mutual information between the current token and the preceding text context grows at least linearly with $t$:
@@ C.1 Visual question answering benchmarks
OK-VQA: OK-VQA(Outside Knowledge VQA)Marino et al. [2019] is a visual question answering dataset that requires external knowledge. The answers to the questions cannot be inferred solely from the image but also need to incorporate common sense or world knowledge. This dataset evaluates the model’s ability in the intersection of vision and knowledge reasoning. We used the acc of validation set to report Our result.
GQA: GQA(Graph Question Answering)Hudson and Manning [2019] generates questions and answers based on image scene graphs, focusing on structured reasoning. It emphasizes logical analysis and challenges the model’s depth of understanding of semantics and context.
VQAV2: VQAv2Goyal et al. [2017] is a new dataset containing open-ended questions about images. These questions require an understanding of vision, language and commonsense knowledge to answer. We used the test split to report Our result.
TextVQA: TextVQASingh et al. [2019] focuses on textual information in images, requiring models to recognize and comprehend text within images to answer questions. It drives research on the integration of visual and textual information, expanding the boundaries of visual question answering.
RealWorldQA: RealWorldQAx.ai [2024] features images and questions sourced from real-world scenarios, encompassing diverse content from daily life. The dataset imposes higher requirements on the model’s generalization ability and adaptability to complex scenes.
DOCVQA: This datasetMathew et al. [2021] comprises over 50,000 questions formulated on more than 12,000 document images. Document Visual Question Answering (DocVQA) aims to promote a ”purpose-driven” perspective within the field of Document Analysis and Recognition, encouraging research that directly addresses practical and semantic understanding of document content.